OpenStudy (anonymous):

I need help with a system of equations problem ! 2x + 4y = 8 2y = x + 8

OpenStudy (anonymous):

Let's use substitution. We can rewrite the second equation as \(x=2y-8\). Substitute that into the first equation for x and we get \(2(2y-8)+4y=8\). We can simplify that to \(4y-16+4y=8\), and then to \(8y=24\), so \(y=3\). If \(y=3\), then \(2x+4(3)=8\), so \(2x=-4\), and \(x=-2\). Test it to make sure our answers are correct, \(2(3)=(-2)+8\), which checks out.